OpenStudy (anonymous):

3 + a + b = a + 3 + b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (austinl):

Commutative Property.

OpenStudy (anonymous):

I believe it's Commutative Property of Addition. I remember this because the numbers can "commute" to other places in the equation and the answers will be the same.

OpenStudy (austinl):

Associative property is, 2(3x) (2*3)x 6x Distributive Property is, a(b + c) = ab + ac 2(3 + 4) = 2×3 + 2×4